5/2 x + 3 =1/2 x +15
Answer by DJ_Sammie(24)

5/2x+3=1/2x+15
you have to get the variable on one side
so subtract 1/2x from both sides
5/2x+3-1/2x=1/2x+15-1/2x
4/2x+3=15
then simplify
2x+3=15
subtract 3 from both sides
2x+3-3=15-3
2x=12
then divide by 2
2x/2=12/2
x=6
x equals 6